OpenStudy (anonymous):

x2+8x+16=0 quadratic equation by factoring

OpenStudy (anonymous):

U can write 8x in such a way that If you add two numbers you will get 8x and if you multiply those two numbers you should get 16x^2... Solving this... U can write 8x as 4x and 4x... If u add these two. U will get 8x and if u multiply u will get 16x^2. got it?

OpenStudy (anonymous):

U will get (x+4)^2=0; So u have two roots, real and equal... x=-4,-4

OpenStudy (tkhunny):

you MUST recognize a Perfect Square Trinomial when you see one. Listen to @waleedimtiaz

OpenStudy (mrnood):

@aj01 (x+2)(x+8) = x^2 +10x +16 so it is not a correct factorisation.